Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 130,000 $ 139,000
Accounts receivable 346,000 487,000
Inventory 567,000 487,000
Plant and equipment, net 817,000 787,000
Investment in Buisson, S.A. 392,000 430,000
Land (undeveloped) 255,000 246,000
Total assets $ 2,507,000 $ 2,576,000
Liabilities and Stockholders' Equity
Accounts payable $ 380,000 $ 339,000
Long-term debt 977,000 977,000
Stockholders' equity 1,150,000 1,260,000
Total liabilities and stockholders' equity $ 2,507,000 $ 2,576,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,324,000
Operating expenses 3,805,120
Net operating income 518,880
Interest and taxes:
Interest expense $ 123,000
Tax expense 208,000 331,000
Net income $ 187,880


The company paid dividends of $77,880 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,507,000-255,000-392,000)+(2,576,000-246,000-430,000)]/2

= 1,880,000

Margin = Net operating income / Sales

= 518,880 / 4,324,000

= 12%

Turnover = Sales /Average operating assets

= 4,324,000 / 1,880,000

= 2.3

ROI =Margin * turnover

= 12*2.3

= 27.60%

Residual income = Net operating income - required return

= 518,880 - (1,880,000*15%)

= 236,880